College Supply Company (CSC) makes three types of drinking glasses: short, medium, and tall. It presently applies overhead using a predetermined rate based on direct labor-hours. A group of company employees recommended that CSC switch to activity-based costing and identified the following activities, cost drivers, estimated costs, and estimated cost driver units for Year 5 for each activity center.

Activity Recommended
Cost Driver
Estimated
Cost
Estimated Cost
Driver Units
Setting up production Number of production runs $ 33,600 120 runs
Processing orders Number of orders 48,000 200 orders
Handling materials Pounds of materials 18,000 9,000 pounds
Using machines Machine-hours 40,000 8,000 hours
Providing quality management Number of inspections 48,000 40 inspections
Packing and shipping Units shipped 40,000 20,000 units
$ 227,600

In addition, management estimated 2,000 direct labor-hours for year 5.

Assume that the following cost driver volumes occurred in February, year 5.

Short Medium Tall
Number of units produced 1,100 500 300
Direct materials costs $ 4,000 $ 2,000 $ 1,500
Direct labor-hours 110 130 120
Number of orders 8 9 5
Number of production runs 3 4 8
Pounds of material 500 800 100
Machine-hours 400 200 300
Number of inspections 2 2 1
Units shipped 1,100 500 200

Direct labor costs were $20 per hour.

Required:

a. Compute a predetermined overhead rate for year 5 for each cost driver recommended by the employees. Also compute a predetermined rate using direct labor-hours as the allocation base.
b. Compute the production costs for each product for February using direct labor-hours as the allocation base and the predetermined rate computed in requirement a.
c. Compute the production costs for each product for February using the cost drivers recommended by the employees and the predetermined rates computed in requirement a. (Note: Do not assume that total overhead applied to products in February will be the same for activity-based costing as it was for the labor-hour-based allocation.)
Answer #1

a)Predetermined overhead rate for each cost driver recommended by employees as below:

Number of production runs = $33600/120=$280 per run
Number of orders = $48000/200= $240 per order
Pound of materials = $18000/9000 = $2 per pound
Machine hours = $40000/8000= $5 per hour
Number of Inspections = $48000/40=$1200 per inspection
Units shipped = $40000/2000 = $2 per unit

predetermined rate using direct labor-hours as the allocation base = $227,600/2000 = $113.8 per hour

b) Solution as below:

Component of cost Short Medium Tall
Material costs (0) 4000 2000 1500
Direct labor hours (i) 110 130 120
Direct labor costs per hour (ii) (given in question) 20 20 20
Overhead labor cost per hour (iii) (computed in part a) 113.8 113.8 113.8
Total cost per hour (iv=ii+iii) 133.8 133.8 133.8
Total cost (v=0+ i*iv) 18718 19394 17556
Number of units (vi) 1100 500 200
Production cost per unit (=v/vi) 17.02 38.79 87.78

c) Solution as below:

Consumption of cost driver units Total Cost
Cost Driver Rate per driver unit ($) Short Medium Tall Short Medium Tall
Number of production runs (i) 280 3 4 8 840 1120 2240
Number of orders (ii) 240 8 9 5 1920 2160 1200
Pounds of materials (iii) 2 500 800 100 1000 1600 200
Machine-hours (iv) 5 400 200 300 2000 1000 1500
Number of inspections (v) 1200 2 2 1 2400 2400 1200
Units shipped (vi) 2 1100 500 200 2200 1000 400
Material cost (vii) 4000 2000 1500
Labor hours (viii) 110 130 120
Direct labor costs (@ $20 per hour) (ix) 2200 2600 2400
Total costs (x = i+ii+iii+iv+v+vi+vii+ix) 16560 13880 10640
Total cost per unit (=x/vi) 15.05 27.76 53.20